6) Predict the sign of Delta S for each of the following changes

K(s) + ½ Br2(s) ? KBr(s)
Answer #1

From standard entropy values

compound/element S(J/mol/K)
K(s) 64.67
Br2(g) 245.3
KBr(s) 95.94

\Delta S =\sum n \times S(products) - \sum n\times S(reactants)

\Delta S =95.94 - [1/2(245.3)+64.67]

\Delta S =95.94 - [187.32]

\Delta S =-91.38\ J/mol/K

Sign of  Delta S = negative